Container A holds a solutions that is 12% acid and container B holds a solution that is 25% acid. How much of each solution should one use to form 7 liters that is 15% acid ?
Answer by ptaylor(2198)

Let x=amount of 12% solution needed
Then 7-x=amount of 25% solution needed
Now we know that the amount of pure acid in the 12% solution (0.12x) plus the amount of pure acid in the 25% solution (0.25(7-x)) has to equal the amount of pure acid after they are mixed together (0.15*7). So, our equation to solve is:
0.12x+0.25(7-x)=0.15*7 get rid of parens
0.12x+1.75-0.25x=1.05 subtract 1.75 from each side
0.12x-0.25x+1.75-1.75=1.05-1.75 collect like terms
-0.13x=-0.70 divide each side by -0.13
x=5.39 liters----------------amount of 12% solution needed
7-x=7-5.39=1.61 liters amount of 25% solution needed
CK
5.39*0.12+1.61*0.25=0.15*7
0.6468+0.4025=1.05
1.049~~~1.05
Hope this helps---ptaylor